# The Nonstationary Newsvendor with (and without) Predictions

### v3

@@ Abstract.
The classic newsvendor model yields an optimal decision for a “newsvendor” selecting a quantity of inventory, under the assumption that the demand is drawn from a known distribution. Motivated by applications such as cloud provisioning and staffing, we consider a setting in which newsvendor-type decisions must be made sequentially, in the face of demand drawn from a stochastic process that is both unknown and nonstationary. All prior work on this problem either (a) assumes that the level of nonstationarity is known, or (b) imposes additional statistical assumptions that enable accurate predictions of the unknown demand.
We study the Nonstationary Newsvendor, with and without predictions. We first, in the setting without predictions, design a policy which we prove (via matching upper and lower bounds) achieves order-optimal regret – ours is the first policy to accomplish this without being given the level of nonstationarity of the underlying demand. We then, for the first time, introduce a model for generic (i.e. with no statistical assumptions) predictions with arbitrary accuracy, and propose a policy that incorporates these predictions without being given their accuracy. We upper bound the regret of this policy, and show that it matches the best achievable regret had the accuracy of the predictions been known. Finally, we empirically validate our new policy with experiments based on three real-world datasets containing thousands of time-series, showing that it succeeds in closing approximately 74% of the gap between the best approaches based on nonstationarity and predictions alone.
@@ 1.3. Literature Review
All the previous studies consider the newsvendor in a static environment where the demand distribution is the same over time. However, in reality the demand distribution is often nonstationary. There are two common practices to resolve this issue. The first is to model the nonstationarity and utilize past demand observations according to the model. One common way is to model the nonstationarity as a Markov chain. For example, [52] applied this idea to inventory management and [5] and [13] applied this idea to revenue management. Another approach is to bound the nonstationarity via a variation budget, which has been applied to stochastic optimization [10], dynamic pricing [28], multi-armed bandit [9], newsvendor problem [27], among others. Some of these works are applicable in the sense that our problem can be mapped to their settings (e.g. multi-armed bandit such as [10, 25, 37, 14]), but these connections do not appear to be fruitful. Finally, as mentioned before, [27] is particularly relevant, so we delay a careful comparison to Sections 2 and 3.
@@ 2. Model: The Nonstationary Newsvendor (without Predictions)
We begin this section with a formal description of the Nonstationary Newsvendor, along with a comparison to the problem of the same name from [27]. Consider a sequence of newsvendor problems over $T$ time periods labeled $t=1,\dots,T$. At the beginning of each time period $t$, the decision-maker selects a quantity $q_{t}\in Q$, where $Q$ is a fixed subset of $\mathbb{R}^{+}$ bounded above by a quantity we denote as $Q_{\max}$. [^2] All of our results carry through if $Q$ is allowed to depend on $t$. Then the period’s demand $d_{t}$ is drawn from an (unknown) demand distribution $D_{t}$, which depends on the time period $t$. Finally a cost is incurred – specifically, there is a (known) per-unit underage cost $b_{t}\in[0,b_{\max}]$ and a (known) per-unit overage cost $h_{t}\in[0,h_{\max}]$, so that the total cost is equal to
@@ 2.1. Demand Distributions
$C(\mu,b,h,q)=\mathbb{E}_{\mathcal{D}_{\mu}}[b(\mathcal{D}_{\mu}-q)^{+}+h(q-\mathcal{D}_{\mu})^{+}].$
@@
    The demand distributions $D_{t}$ in [27] are assumed to be of the form $D_{t}=\mu_{t}+\epsilon_{t}$, where $\mu_{t}$ is the mean of $D_{t}$ that drifts across time and $\epsilon_{t}$ is the noise distribution that is i.i.d., continuous, and bounded. Effectively, the demand distribution preserves its “shape” at all times. We make no such restriction.
@@ 2.2. Demand Variation
(2.2) $V_{\bm{\mu}}=\max_{\{t_{0},\dots,t_{K}\}\in\mathcal{P}}\left\{\sum_{k=1}^{K}\left|\mu_{t_{k}}-\mu_{t_{k-1}}\right|^{2}\right\},$
@@ 2.3. Performance Metric: Regret
where the expectation is with respect to the randomness of the realized demands. In words, the regret measures the difference between the (expected) total cost incurred by the policy and that of a clairvoyant that knows the underlying demand distributions $\bm{D}=\{D_{1},\dots,D_{T}$}. [^10] Note that this is different from a clairvoyant that knows the realized demands $d_{1},\ldots,d_{T}$. Such a clairvoyant would incur zero cost.
@@ 3. Solution to the Nonstationary Newsvendor (without Predictions)
For any variation parameter $v\in[0,1]$, and any policy $\pi$ (which may depend on $v$), we have
@@
    As an aside, this discussion raises a natural question: is the disconnect here between stationary (i.i.d.) demand and variation parameter $v=0$ a consequence of our use quadratic variation, and would the same disconnect arise for other measures of demand variation? In Appendix E, we answer both questions in the affirmative by showing that if the exponent $2$ in the demand variation (Eq. 2.2) is instead some $q\geq 0$, then Proposition 1 generalizes to a lower bound of $\Omega(T^{(1+q+v)/(2+q)})$. Thus, for any $q>0$, the case of variation parameter $v=0$ is meaningfully more general than stationary demand.
@@
    Continuous Newsvendor: A similar “story” plays out in the setting of [27], which recall requires the additional assumption that both the demand distributions and the possible order quantities be continuous. [27] show an optimal achievable regret of $\Theta(T^{(1+v)/2})$, which can be contrasted with the stationary (i.i.d.) setting for which an $\Omega(\log T)$ lower bound exists ([11]). The following table summarizes these lower bounds:
@@ 4.1. Model
    For any policy $\pi$ (which may depend on $a$) that is solely a function of the predictions (i.e. does not depend on the observed demands), we have
@@ 4.2. Extreme Cases
Fix any variation parameter $v\in[0,1]$ and any accuracy parameter $a\in[0,1]$. For any policy $\pi$ (which may depend on $v$ and $a$), we have
@@
For any policy that does not depend on $v$ or $a$, there exists a problem instance such that $a\neq(3+v)/4$, and the policy incurs regret at least $cT^{\max\{(3+v)/4,a\}}$ on the instance, where $c>0$ is a universal constant.
Proposition 3 states that the best we can hope for, when $v$ and $a$ are unknown, is a worst-case regret of at least $\Omega(T^{\max\{(3+v)/4,a\}})$. Indeed, it implies that no algorithm can achieve regret $O(T^{f(v,a)})$ for a function $f:[0,1]\times[0,1]\to[0,1]$ satisfying $f(v,a)\leq\max\{(3+v)/4,a\}$ for all $v,a\in[0,1]$ and $f(v,a)<\max\{(3+v)/4,a\}$ for some $v,a\in[0,1]$. This lower bound is easily achieved, for example by applying the Shrinking-Time-Window Policy or the Prediction Policy (or any blind randomization of the two). The proof of Proposition 3 is in Appendix F. In contrast to Case 1, the lower bound construction here relies heavily on the fact that we do not know which one of $(3+v)/4$ and $a$ is smaller.
@@ 4.3. Final Solution
Theorem 2 (Upper Bound: Known $v$ and Unknown $a$).
$\mathcal{R}^{\pi^{\mathrm{PERP}}}(T)\leq CT^{\min\{(3+v)/4,a\}}\sqrt{\log T},$
where $C$ is a universal constant.
@@ 5.1. Experiments on Synthetic Data
Figure 2. The costs of NO-PRED, PURE-PRED, and PERP when (a) the variation parameter $v$ is fixed, and (b) the accuracy parameter $a$ is fixed. Each dot represents the cost of the corresponding policy on a given instance.
For each demand sequence and corresponding prediction, we ran NO-PRED, PURE-PRED, and PERP with equal overage and underage costs, and scaling constants $\kappa=\gamma=1$. The variation parameter $v$ in PERP was calculated using the past demands of the pre-fixed 30 time periods by the definition in Section 2.2. We calculated the parameters $v$ and $a$ by their definitions (given in Section 2.2 and Section 4.1, respectively), scaled appropriately to make them lie in $[0,1]$. The resulted scatter plots are shown in Fig. 2. In (a), $v$ is fixed, so the cost of NO-PRED (blue dots) is approximately the same for all instances. The cost of PURE-PRED (orange dots) is approximately exponential in $a$, which follows by 1a). In (b), $a$ is fixed, so the cost of PURE-PRED is approximately the same for all instances. The cost of NO-PRED is approximately exponential in $v$, which follows by Theorem 1. Note that in both (a) and (b), the cost of PERP (green dots) is close to the minimal cost of NO-PRED and PURE-PRED across all instances, showing that PERP’s performance is robust in both $v$ and $a$.
@@ Appendix E General Variation
where $\mathcal{P}$ is the set of all partitions. The choice of quadratic variation follows from previous literature ([28, 27]). We can also define the demand variation using what we will call $q$-variation for some $0\leq q<\infty$:
$V_{\bm{\mu}}=\max_{\{t_{0},\dots,t_{K}\}\in\mathcal{P}}\left\{\sum_{k=1}^{K}\left|\mu_{t_{k}}-\mu_{t_{k-1}}\right|^{q}\right\}.$
In the special case of $q=1$, this is the total variation, which has also been used extensively in previous literature ([10, 25, 37, 14]). We prove the following lower bound, which generalizes Proposition 1.
Proposition 4 (Lower Bound: Nonstationary Newsvendor with $q$-variation).
Suppose the demand variation is defined using $q$-variation and $V_{\bm{\mu}}\leq T^{v}$. For any variation parameter $v\in[0,1]$, and any policy $\pi$ (which may depend on $v$), we have
$\mathcal{R}^{\pi}(T)\geq cT^{(1+q+v)/(2+q)},$
@@
Similar to the proof of Proposition 1, we construct the following worst-case problem instance: divide the time horizon $T$ into cycles of length $T^{(2-2v)/(2+q)}$ (since the analysis below is compatible with scaling, for simplicity we assume $T^{(2-2v)/(2+q)}$ is an integer), so there are $T^{(q+2v)/(2+q)}$ cycles. Assume that the demand distribution within each cycle is a Bernoulli distribution that equals to $1$ with probability $p$ and equals to 00 with probability $1-p$. At the beginning of each cycle, we set the $p$ of the upcoming cycle to be either $\frac{1}{2}+\frac{1}{\sqrt{20}}T^{(v-1)/(2+q)}$ or $\frac{1}{2}-\frac{1}{\sqrt{20}}T^{(v-1)/(2+q)}$, each with probability $\frac{1}{2}$. Set $Q=\mathbb{R}_{+}$ and $b_{t}=h_{t}=1$ for all $t$, so the optimal ordering amount is the median of the demand distribution at each time.
First we show that the demand variation $V_{\bm{\mu}}$ is at most $T^{v}$. Note that since $\mu_{t}=p$ is fixed within each cycle, only the times between cycles contribute to the demand variation. Suppose the cycle changes between time $t$ and $t+1$, then $\left(\mu_{t+1}-\mu_{t}\right)^{q}\leq(\frac{2}{\sqrt{20}}T^{(v-1)/(2+q)})^{q}\leq T^{q(v-1)/(2+q)}$. Because there are $T^{(q+2v)/(2+q)}$ cycles, $V_{\bm{\mu}}\leq T^{(q+2v)/(2+q)}\cdot T^{q(v-1)/(2+q)}=T^{v}$.
Then, following the calculations in the proof of Proposition 2, the probability of making the wrong guess of $p$ in the current cycle is at least $1-\sqrt{\frac{13}{40}}$. Also, each wrong guess of $p$ incurs a regret of $\frac{1}{\sqrt{5}}T^{(v-1)/(2+q)}$ at each time period. Therefore over $T$ time periods the expected total regret of any General Policy $\pi$ satisfies
$\mathcal{R}^{\pi}(T)$ $\geq\left(1-\sqrt{\frac{13}{40}}\right)\frac{1}{\sqrt{5}}T^{(v-1)/(2+q)}\cdot T=\left(1-\sqrt{\frac{13}{40}}\right)\frac{1}{\sqrt{5}}T^{(1+q+v)/(2+q)}.$

### v4

@@ Abstract.
Problem definition: The classic newsvendor model yields an optimal decision for a “newsvendor” selecting a quantity of inventory, under the assumption that the demand is drawn from a known distribution. Motivated by applications such as cloud provisioning and staffing, we consider a setting in which newsvendor-type decisions must be made sequentially, in the face of demand drawn from a stochastic process that is both unknown and nonstationary. All prior work on this problem either (a) assumes that the level of nonstationarity is known, or (b) imposes additional statistical assumptions that enable accurate predictions of the unknown demand. Our research tackles the Nonstationary Newsvendor without these assumptions, both with and without predictions.
Methodology/results: We first, in the setting without predictions, design a policy which we prove (via matching upper and lower bounds) achieves order-optimal regret – ours is the first policy to accomplish this without being given the level of nonstationarity of the underlying demand. We then, for the first time, introduce a model for generic (i.e. with no statistical assumptions) predictions with arbitrary accuracy, and propose a policy that incorporates these predictions without being given their accuracy. We upper bound the regret of this policy, and show that it matches the best achievable regret had the accuracy of the predictions been known.
Managerial implications: Our findings provide valuable insights on inventory management. Managers can make more informed and effective decisions in dynamic environments, reducing costs and enhancing service levels despite uncertain demand patterns. This study advances understanding of sequential decision-making under uncertainty, offering robust methodologies for practical applications with nonstationary demand. We empirically validate our new policy with experiments based on three real-world datasets containing thousands of time-series, showing that it succeeds in closing approximately 74% of the gap between the best approaches based on nonstationarity and predictions alone.
This paper is published at https://pubsonline.informs.org/doi/10.1287/msom.2024.1168.
@@ 1.3. Literature Review
All the previous studies consider the newsvendor in a static environment where the demand distribution is the same over time. However, in reality the demand distribution is often nonstationary. There are two common practices to resolve this issue. The first is to model the nonstationarity and utilize past demand observations according to the model. One common way is to model the nonstationarity as a Markov chain. For example, [52] applied this idea to inventory management and [5] and [13] applied this idea to revenue management. Another approach is to bound the nonstationarity via a variation budget, which has been applied to stochastic optimization [10], dynamic pricing [28], multi-armed bandit [9], newsvendor problem [27], among others. Some of these works are applicable in the sense that our problem can be mapped to their settings (e.g. multi-armed bandit such as [10, 25, 37, 14]), but these connections do not appear to be fruitful. In particular, the multi-armed bandit papers cited above typically consider a limited-feedback setting rather than the full-feedback setting explored in this work. Related to feedback, while our study provides a complete characterization of the regret behavior for the nonstationary newsvendor problem with uncensored demand, practical applications often involve censored demand. The nonstationary newsvendor problem under censored demand is an interesting direction for future research.
Beyond the bandit literature, it is worth mentioning recent work on online convex optimization (OCO) with limited nonstationarity. When the level of nonstationarity is known, the standard first-order OCO algorithms can be modified with carefully chosen restarts and updating rules ([10],[59], [57]). There are also recent works that concern unknown nonstationarity, such as [60], [56], [61], and [58]. Finally, as mentioned before, [27] is particularly relevant, so we delay a careful comparison to Sections 2 and 3.
@@ 2. Model: The Nonstationary Newsvendor (without Predictions)
We begin this section with a formal description of the Nonstationary Newsvendor, along with a comparison to the problem of the same name from [27]. Consider a sequence of newsvendor problems over $T$ time periods labeled $t=1,\dots,T$. At the beginning of each time period $t$, the decision-maker selects a quantity $q_{t}\in Q$, where $Q$ is a fixed subset of $\mathbb{R}^{+}$ bounded above by a quantity we denote as $Q_{\max}$. [^2] All of our results carry through if $Q$ is allowed to depend on $t$. Then the period’s demand $d_{t}$ is drawn from an (unknown) demand distribution $D_{t}$, which depends on the time period $t$. These demand distributions are independent over time. Finally a cost is incurred – specifically, there is a (known) per-unit underage cost $b_{t}\in[0,b_{\max}]$ and a (known) per-unit overage cost $h_{t}\in[0,h_{\max}]$, so that the total cost is equal to
@@ 2.1. Demand Distributions
$C(\mu,b,h,q)=\mathbb{E}_{d\sim\mathcal{D}_{\mu}}[b(d-q)^{+}+h(q-d)^{+}].$
@@
    The demand distributions $D_{t}$ in [27] are assumed to be of the form $D_{t}=\mu_{t}+\epsilon_{t}$, where $\mu_{t}$ is the mean of $D_{t}$ that drifts across time and $\epsilon_{t}$ is the noise distribution that is i.i.d., continuous, and bounded. Effectively, the demand distribution fall into a non-parametric family of distributions with the same “shape”. In contrast, our demand distributions fall into a parametric family of distributions, though not necessarily of the same “shape”.
@@ 2.2. Demand Variation
(2.2) $V_{\bm{\mu}}=\max_{0\leq K\leq T-1}\max_{\{t_{0},\dots,t_{K}\}\in\mathcal{P}}\left\{\sum_{k=1}^{K}\left|\mu_{t_{k}}-\mu_{t_{k-1}}\right|^{2}\right\},$
@@ 2.3. Performance Metric: Regret
where the expectation is with respect to the randomness of the realized demands. Recall that the demand distributions are independent, so $q_{t}^{*}$ as defined in Eq. (2.1) depends only on $D_{t}$. In words, the regret measures the difference between the (expected) total cost incurred by the policy and that of a clairvoyant that knows the underlying demand distributions $\bm{D}=\{D_{1},\dots,D_{T}$}. [^10] Note that this is different from a clairvoyant that knows the realized demands $d_{1},\ldots,d_{T}$. Such a clairvoyant would incur zero cost.
@@ 3. Solution to the Nonstationary Newsvendor (without Predictions)
For any variation parameter $v\in[0,1]$, and any policy $\pi$ (which may depend on the knowledge of $v$), we have
@@
    As an aside, this discussion raises a natural question: is the disconnect here between stationary (i.i.d.) demand and variation parameter $v=0$ a consequence of our use quadratic variation, and would the same disconnect arise for other measures of demand variation? In Appendix E, we answer both questions in the affirmative by showing that if the exponent $2$ in the demand variation (Eq. 2.2) is instead some ${\theta}\geq 0$, then Proposition 1 generalizes to a lower bound of $\Omega(T^{(1+{\theta}+v)/(2+{\theta})})$. Thus, for any ${\theta}>0$, the case of variation parameter $v=0$ is meaningfully more general than stationary demand.
@@
    Continuous Newsvendor: A similar “story” plays out in the setting of [27], which recall (among other key differences with our model, as described in Section 2.1) requires the additional assumption that both the demand distributions and the possible order quantities be continuous. [27] show an optimal achievable regret of $\Theta(T^{(1+v)/2})$, which can be contrasted with the stationary (i.i.d.) setting for which an $\Omega(\log T)$ lower bound exists ([11]). The following table summarizes these lower bounds:
@@ 4.1. Model
    For any policy $\pi$ (which may depend on the knowledge of $a$) that is solely a function of the predictions (i.e. does not depend on the observed demands), we have
@@ 4.2. Extreme Cases
Fix any variation parameter $v\in[0,1]$ and any accuracy parameter $a\in[0,1]$. For any policy $\pi$ (which may depend on the knowledge of $v$ and $a$), we have
@@
For any policy that does not depend on the knowledge of $v$ or $a$, there exists a problem instance such that $a\neq(3+v)/4$, and the policy incurs regret at least $cT^{\max\{(3+v)/4,a\}}$ on the instance, where $c>0$ is a universal constant.
Proposition 3 states that the best we can hope for, when $v$ and $a$ are unknown, is a worst-case regret of at least $\Omega(T^{\max\{(3+v)/4,a\}})$. Indeed, it implies that no algorithm can achieve regret $O(T^{f(v,a)})$ for a function $f:[0,1]\times[0,1]\to[0,1]$ satisfying $f(v,a)\leq\max\{(3+v)/4,a\}$ for all $v,a\in[0,1]$ and $f(v,a)<\max\{(3+v)/4,a\}$ for some $v,a\in[0,1]$. Note that Proposition 3 shows there exists a pair of $v$ and $a$, and a corresponding problem instance, such that this lower bound holds. This is in contrast to a result showing that for any pair of $v$ and $a$, there exists a problem instance such that the lower bound holds, as is common in the literature (e.g. Theorem 1 of [28]). This lower bound is easily achieved, for example by applying the Shrinking-Time-Window Policy or the Prediction Policy (or any blind randomization of the two). The proof of Proposition 3 is in Appendix F. In contrast to Case 1, the lower bound construction here relies heavily on the fact that we do not know which one of $(3+v)/4$ and $a$ is smaller.
@@ 5.1. Experiments on Synthetic Data
Figure 2. The costs of NO-PRED, PURE-PRED, PERP, and OPT when (a) the variation parameter $v$ is fixed, and (b) the accuracy parameter $a$ is fixed. Each dot represents the cost of the corresponding policy on a given instance.
For each demand sequence and corresponding prediction, we ran NO-PRED, PURE-PRED, and PERP with equal overage and underage costs, and scaling constants $\kappa=\gamma=1$. Because the experiment was synthetic, the true underlying demand distribution was known at each time period. Therefore we also ran OPT as a benchmark, which simply ordered the optimal quantile at each time period. The variation parameter $v$ in PERP was calculated using the past demands of the pre-fixed 30 time periods by the definition in Section 2.2. We calculated the parameters $v$ and $a$ by their definitions (given in Section 2.2 and Section 4.1, respectively), scaled appropriately to make them lie in $[0,1]$. The resulted scatter plots are shown in Fig. 2. In (a), $v$ is fixed, so the cost of NO-PRED (blue dots) is approximately the same for all instances. The cost of PURE-PRED (orange dots) is approximately exponential in $a$, which follows by 1a). In (b), $a$ is fixed, so the cost of PURE-PRED is approximately the same for all instances. The cost of NO-PRED is approximately exponential in $v$, which follows by Theorem 1. Note that in both (a) and (b), the cost of PERP (green dots) is close to the minimal cost of NO-PRED and PURE-PRED across all instances, showing that PERP’s performance is robust in both $v$ and $a$.
@@ Appendix E General Variation
where $\mathcal{P}$ is the set of all partitions. The choice of quadratic variation follows from previous literature [28, 27]. We can also define the demand variation using what we will call $\theta$-variation for some $0\leq{\theta}<\infty$:
$V_{\bm{\mu}}=\max_{\{t_{0},\dots,t_{K}\}\in\mathcal{P}}\left\{\sum_{k=1}^{K}\left|\mu_{t_{k}}-\mu_{t_{k-1}}\right|^{\theta}\right\}.$
In the special case of ${\theta}=1$, this is the total variation, which has also been used extensively in previous literature [10, 25, 37, 14]. We prove the following lower bound, which generalizes Proposition 1.
Proposition 4 (Lower Bound: Nonstationary Newsvendor with $\theta$-variation).
Suppose the demand variation is defined using ${\theta}$-variation and $V_{\bm{\mu}}\leq T^{v}$. For any variation parameter $v\in[0,1]$, and any policy $\pi$ (which may depend on the knowledge of $v$), we have
$\mathcal{R}^{\pi}(T)\geq cT^{(1+\theta+v)/(2+\theta)},$
@@
Similar to the proof of Proposition 1, we construct the following worst-case problem instance: divide the time horizon $T$ into cycles of length $T^{(2-2v)/(2+{\theta})}$ (since the analysis below is compatible with scaling, for simplicity we assume $T^{(2-2v)/(2+{\theta})}$ is an integer), so there are $T^{(\theta+2v)/(2+\theta)}$ cycles. Assume that the demand distribution within each cycle is a Bernoulli distribution that equals to $1$ with probability $p$ and equals to 00 with probability $1-p$. At the beginning of each cycle, we set the $p$ of the upcoming cycle to be either $\frac{1}{2}+\frac{1}{\sqrt{20}}T^{(v-1)/(2+\theta)}$ or $\frac{1}{2}-\frac{1}{\sqrt{20}}T^{(v-1)/(2+\theta)}$, each with probability $\frac{1}{2}$. Set $Q=\mathbb{R}_{+}$ and $b_{t}=h_{t}=1$ for all $t$, so the optimal ordering amount is the median of the demand distribution at each time.
First we show that the demand variation $V_{\bm{\mu}}$ is at most $T^{v}$. Note that since $\mu_{t}=p$ is fixed within each cycle, only the times between cycles contribute to the demand variation. Suppose the cycle changes between time $t$ and $t+1$, then $\left(\mu_{t+1}-\mu_{t}\right)^{\theta}\leq(\frac{2}{\sqrt{20}}T^{(v-1)/(2+\theta)})^{\theta}\leq T^{\theta(v-1)/(2+\theta)}$. Because there are $T^{(\theta+2v)/(2+\theta)}$ cycles, $V_{\bm{\mu}}\leq T^{(\theta+2v)/(2+\theta)}\cdot T^{\theta(v-1)/(2+\theta)}=T^{v}$.
Then, following the calculations in the proof of Proposition 2, the probability of making the wrong guess of $p$ in the current cycle is at least $1-\sqrt{\frac{13}{40}}$. Also, each wrong guess of $p$ incurs a regret of $\frac{1}{\sqrt{5}}T^{(v-1)/(2+\theta)}$ at each time period. Therefore over $T$ time periods the expected total regret of any General Policy $\pi$ satisfies
$\mathcal{R}^{\pi}(T)$ $\geq\left(1-\sqrt{\frac{13}{40}}\right)\frac{1}{\sqrt{5}}T^{(v-1)/(2+\theta)}\cdot T=\left(1-\sqrt{\frac{13}{40}}\right)\frac{1}{\sqrt{5}}T^{(1+\theta+v)/(2+\theta)}.$